Gig Harbor Boating is the wholesale distributor of a small recreational catamaran sailboat. Management has prepared the following summary data to use in its annual budgeting process:

Budgeted unit sales 520
Selling price per unit $ 1,980
Cost per unit $ 1,660
Variable selling and administrative expense (per unit) $ 60
Fixed selling and administrative expense (per year) $ 105,000
Interest expense for the year $ 14,000

Required:

Prepare the company’s budgeted income statement for the year.
Answer #1

Solution

Gig Harbor Boating
Budgeted Income Statement
Sales revenue $ 1,029,600.00
Cost of goods sold $      863,200.00
Gross profit $      166,400.00
Selling and Administrative expenses $      136,200.00
Operating income   $        30,200.00
Interest expense $        14,000.00
Net income $        16,200.00

Working

Gig Harbor Boating
Budgeted Income Statement
Sales revenue =520*1980
Cost of goods sold =520*1660
Selling and Administrative expenses =(60*520)+105000